# Tightrope-walking rationality in action: Feyerabendian insights for the foundations of quantum mechanics

### v2

@@ 1 Introduction
In the area of foundations of physics, the analysis of technical aspects of physical theories and experimental results has the purpose of elucidating the conceptual pillars of the theories, which we use, in turn, to interpret what we see in experiments. In some cases, work on these foundations impact only our understanding of theories. Thus, it is valuable but not ground-breaking; it is interesting for scholars, but without much broader implications. In other cases, it shakes the whole edifice of theoretical physics, and it has important implications for fundamental issues in philosophy of science and philosophy more generally. Thus, it is interesting for a broader set of philosophers, alongside physicists. This was certainly the case, and it still is [^1] It is the case also for the conceptual aspects of research on quantum gravity., for the foundations of quantum mechanics, on whose puzzling aspects many prominent philosophers have exercised their minds from its very first developments.
@@
In the following we focus on the contribution by P. K. Feyerabend (PKF) to the debate on the foundations and interpretation of quantum mechanics, on which he maintained a keen interest from early on but to which he contributed most actively in the decades 1950s-1960s. The most relevant works will be those collected in the first volume of PKF’s papers [1], to which we will refer throughout (rather than referring to the original publications reprinted in it).
@@
We will discuss a number of interesting points made by PFK. We group them in three categories. The first is about the general relation between physics and philosophy, which serves also as a framework for the philosophical analysis of quantum mechanics (and philosophy of physics at large). The second concerns the debate between realism and instrumentalism, applied to the interpretation of quantum mechanics, but also, again, as two broader perspectives on scientific theories. The third deals with specific insights provided by PKF on quantum mechanics itself, concerning the interpretation of quantum states, probabilities, Bohr’s view and hidden variables.
@@
A result of our scan, beyond the specific insights on quantum mechanics, is to illustrate how PFK constantly aimed to achieve a clearer understanding of the various issues he was facing, finding compelling rational arguments in favor of one particular viewpoint, while at the same time avoiding any form of dogmatism, even regarding the rational arguments he had just put forward, and regarding rationality itself (or the criteria adopted for evaluating his own arguments). This struggle characterizes his own version of rational analysis, dubbed ‘tightrope-walking rationality’ by Farrell [2], which we then see in action, in the context of the foundations of quantum mechanics and related philosophical issues. The dogmatic version of the two sides of any argument, of the two positions being confronted (here in particular realism vs instrumentalism will be central), is what we see on the right and on the left of the tightrope we are forced to walk on, and we should avoid falling toward either side. The need to avoid dogmatism of any sort appears to be the strongest driver in PKF’s analysis, to the point that any perspective that is at risk of turning into a dogmatic position or facilitating a dogmatic attitude is immediately counter-argued exactly on that basis. Thus, avoiding dogmatism turns quickly from a normative methodological rule for philosophical inquiry into an evaluative criterion of philosophical positions in itself. We will emphasize various concrete instances of application of such tightrope-walking rationality in what follows.
@@
Before we start, it is worth emphasizing what this contribution is about and what is not aiming at. It is not intended as a scholarly contribution about PKF’s work and thought. We do not attempt a precise or detailed analysis of what PKF really meant or of how PKF’s thinking evolved over the years. We do not frame PKF’s thinking within the historical debate on quantum foundations, and we do not assess his actual impact on the same debate. For proper scholarly work on PKF, especially in relation to contemporary debates in philosophy of science (e.g. on realism, perspectivism, pluralism, scientific methodology, etc), see for example [3, 4, 5] or the cited [2]. What we aim to do is much more limited. Being interested ourselves in the many puzzles of quantum foundations, we scan PKF’s writings for useful insights that could help us improving our own understanding of them, and could suggest ways to make progress, technical or conceptual. In particular, we ask PKF for advice about how one should seek some form of scientific realism behind quantum mechanics, about the complex relation between a scientific theory and reality, and about how to be critical toward our own attitudes when presented with a new set of phenomena and a new candidate description of them. Also in this respect, our contribution should be seen as a modest addition to the scholarly work on this topic, examples being [6, 7] It will be, then, an exercise in picking up useful inputs for current research, and suggestive stimuli to be put to fruition at a second stage. We will try, of course, to be careful in our reading and understanding PKF’s thoughts, and to avoid misinterpretations or cherry-picking or other forms of unfaithfulness to PKF’s intentions and ideas. In case we fail despite our best efforts, and some cherry-picking turns out to be happening, even if unintended, we hope that at least the cherries are good.
@@ 2 Relation between physics and philosophy, in the foundations of QM
An immediate point is PKF’s preference for philosophical analyses that, when not directly done by physicists, it is proceeding in parallel with scientific developments and to some extent guided by them. On the one hand, an important goal of the philosophy of physics is, in PKF’s view, to contribute to the scientific progress; on the other hand, a physical argument or result that supports or suggests a certain philosophical perspective becomes the best argument in its favor and has more weight than abstract philosophical reasoning going in the same or opposite direction. This applies to metaphysical issues, as well as epistemological or methodological ones. In this sense one can understand the many occasions in which PKF expresses discomfort or outright annoyance with philosophical positions that appear oblivious to the hard-won results by scientists or that are based on a superficial picture of scientific theories, and of how they have been obtained and tested, or with methodological rules that have no relation with the practice of science.
We would suggest that a way to understand the spirit of PKF’s attitude in this regard is to consider his tightrope-walking rationality not so much as an abstract theory of rationality to be then applied (or imposed) onto science and scientists, but quite the contrary, as a model of rationality deduced or abstracted from the way actual scientists work, from the necessarily fallibilist and wavering proceeding of real-world science.
In the context of a general survey on scientific realism in physics, for example, PKF praises the ideas of Maxwell and Mach because “They are also more subtle. They were developed in close connection with research […]. But one feels a sense of relief when transferred from the fruitless technicalities and ontological primitivisms of modern ’philosophers’ to the brief, simple, but profound remarks of these scientists”. [[1], page 11].
@@
Lakatos says I use ’some inconsistencies and waverings in Bohr’s position for a crude apologetic falsification of Bohr’s philosophy’. This is the language of the rationalist who looks at research from the safe distance of his office and wants to run science according to one method and base it on one philosophy and who fails to understand that researchers adapt their ideas to the always changing research situation. Bohr does not ’waver’ because he does not know what to do, but because new discoveries are constantly being made and because he takes them into account. - [1], page 294, footnote 100
@@
Ultimately, physical arguments (and results) trump philosophical prejudices, being methodological or ontological. This is as it should be, to make science progress. On the methodological side, this attitude is also at the basis of PKF’s methodological anarchism [8], since it allows for different methodological prescriptions to be employed in different circumstances, and any given methodological rule to be violated, if deemed useful to achieve scientific progress.
@@
Does this mean that philosophy plays no role in scientific progress? Obviously not. Philosophical arguments can guide or support physics, a certain philosophical perspective can prove fertile in driving scientific progress or solving conceptual puzzles encountered by scientists. Ontological preferences are means to physical understanding; they are either starting points for new scientific constructions or results of (analysis of) existing ones. Philosophical arguments can also be a good motivation to oppose a scientific theory or viewpoint, but only as starting point to develop alternative scientific theories. It is at the level of physics that the debate will have to be settled, with the physical reasons being given appropriate (and greater) weight. Philosophical arguments can win a philosophical battle, never one about physics. This implies, in particular, that philosophical as well as physical arguments should be considered in relation to specific theories, those that we are scrutinizing for viability, and not in general, abstract terms.
@@
PKF is very clear on this point, when discussing the debate around the heliocentric view and then using it to elucidate the parallel debate on the realist vs instrumentalist interpretation of quantum theory:
”[…] There exists very impressive physical arguments against a realistic interpretation of quantum theory. There are also more philosophical arguments trying to establish the same result, viz. that the quantum theory is an instrument of prediction from which no realistic consequences can be drawn. These philosophical arguments proceed from the assumption that only observational terms are candidates for a realistic interpretation, an assumption which can be refuted once and for all by philosophical reasoning. Thinkers which are conversant with philosophy only will assume that this settles the matter — which is far from being true. - [1], page 185-186
@@
This is a warning that should be raised also in front of physicists trying to wrap their theories in fancy philosophical clothes, trying to make their insights look deeper, or trying to call philosophical frameworks in further support of their physical theories. This move can have value but one has to be very careful. Bad philosophical arguments can end up obscuring good physical motivations and bring unnecessary trouble to a physical theory or viewpoint that could have stood on more solid grounds without. This is what has happened, according to PKF, to the ”instrumentalist”  interpretation of quantum mechanics. In the attempt to further justify such position beyond the physical arguments by linking it with more general positivism, quantum physicists ended up offering easy attacking points to their opponents. Attacking their positivist position and refuting it with good philosophical arguments made their view of quantum theory appear weaker, despite their being in possession of strong physical results supporting it. So strong, in fact, that they did not need to be further justified in terms of any positivist philosophy. In the same spirit, PKF warns about expressing or interpreting physical theories in a language adapted to or coming from a specific philosophical perspective. In the specific case of quantum transitions, for example, PKF insists that the absence of intermediate states is something to be argued for because of the empirical evidence and of the content of quantum theory that successfully accounts for it, while arguing for this conclusion or justifying the theory on the basis of a philosophical doctrine (here, positivism) does not help:
@@
[…] it should be noted that the proposed solution does not contain any reference to knowledge, or observability. It is not asserted that during the time of transfer A and B may be in some state which is unknown to us, or which cannot be observed. For the quantum postulate does not merely exclude the knowledge of, or the observability of, the intermediate states; it excludes these intermediate states themselves. Nor must the argument be read as asserting, as is implied in many presentations by positivist-minded physicists, that the intermediate states do not exist because they cannot be observed. For it refers to a postulate (the quantum postulate) which deals with existence, and not with observability. It is at this point that the most misleading arguments occur. Physicists who have adopted the positivistic principle that things which cannot be observed do not exist try to justify the indefiniteness of state descriptions by a combined reference to the fact that they cannot be observed and to this principle. Philosophers immediately expose the fallacy of the argument (if they are anti-positivists, that is) and think that they have thereby shown the existence, or at least the physical possibility, of sharp states. This is, of course, not correct, for from the fact that a certain argument has been found to be fallacious it does not follow that a better argument does not exist. But this better argument is hardly ever used by the physicists, which creates the impression that positivism is indeed the only source of the peculiar features of the present quantum theory. The emphasis upon the absence of predictability is not satisfactory either. For this way of speaking would again suggest that we could perhaps predict better if we only knew more about the things that exist in the universe, whereas Bohr’s suggestion denies that there are things whose detection would make our knowledge more definite. […] [1], page 187-190
@@
[…] modern physics has found new physical reasons why its own most important theory, the quantum theory, cannot be anything but an instrument of prediction. These reasons are of precisely the same character as were Ptolemy’s: a realistic interpretation of the quantum theory is bound to lead to incorrect predictions. Admittedly, in the arguments usually presented these physical difficulties are almost buried beneath an unacceptable positivism. However, this does not mean that they do not exist and that no ’further shot had been fired’ since the time of Bellarmine and Berkeley.” - [1], page 187-190
@@
”the instrumentalism in the quantum theory is not a purely philosophical affair that can be disputed away by general arguments in favour of realism. The argument ’Quantum theoretical instrumentalism is a result of positivism; positivism is false; hence, we must interpret the quantum theory in a realistic fashion’, is completely irrelevant and also very misleading. It is very misleading because it suggests […] that the reason why it was not so interpreted was only philosophical prejudice. And it is irrelevant because it does not proceed a single step on the way to resolving the physical difficulties which are connected with the realistic position in microphysics. A realistic alternative to the idea of complementarity is likely to be successful only if it implies that certain experimental results are not strictly valid. It therefore demands the construction of a new theory, as well as demonstration that this new theory is experimentally at least as valuable as the theory that is being used at present. This is a formidable task indeed, and a task that is not even recognized by the purely philosophical champions of realism in microphysics. - [1], page 193-194
@@
I consider Bohm’s book to be a major contribution to the contemporary philosophy of nature I must at once add that there are many things in it which I cannot accept and that more especially his discussion of the problem of induction seems to me to be highly unsatisfactory. Bohm’s physical ideas are original, refreshing, and sorely needed in a time of complacency with respect to fundamentals. But the philosophical standpoint taken up with respect to both physics and cosmology is traditional, and perhaps even reactionary - [1], page 220
In the same discussion PKF introduces another important element: a sound position can become quickly counterproductive and represent an obstacle to further progress the very moment it become dogmatic, no matter how many solid arguments support it. A position that is well supported by physical arguments and observations, that is embodied in the best physical framework we were able to find, and that it is proving fertile for further scientific progress becomes dogmatic when it is declared to be necessarily part of any future theory, to stand without possible alternatives, rather than merely current ones, as something having absolute validity. Also here one can see the ideal tightrope-walking philosopher or scientist moving the balance quickly away from the full (and possibly uncritical) embracing of a theory and towards a more skeptical or simply actively critical attitude, to avoid falling into the pit of dogmatism. This dogmatic turn is what PKF attributes to the ”orthodox” Copenhagen interpretation and to the physicists holding it. In his words:
The idea of complementarity can be interpreted in two different ways. It can be interpreted as an attempt to provide an intuitive picture for an existing theory, viz. wave mechanics, and as a heuristic principle guiding future research. This interpretation is undogmatic as it admits the possibility of alternatives, and even of preferable alternatives. A physicist who looks at complementarity in this way will regard it as an interesting fact about quantum theory that it is compatible with a relational point of view where interaction is a necessary condition of the meaningful applicability of terms which within classical physics (relativity included) are definable without such reference. He will also point out that there exist no satisfactory alternatives. But he will never go as far as to assert that such alternatives will never be found, or that they would be logically inconsistent, or that they would contradict the facts. But Bohr’s idea of complementarity can also be interpreted in a different way. It can be interpreted as a basic philosophical principle which is incapable of refutation and to which any future theory must conform. Bohr himself most certainly took this stronger point of view. - [1], page 221-222
@@
it has been attempted, both by Bohr, and by some other members of the Copenhagen circle, to give greater credibility to these ideas by incorporating them into a whole philosophical (ontological) system that comprises physics, biology, psychology, sociology and perhaps even ethics. Now the attempt to relate physical ideas to a more general background and the correlated attempt to make them intuitively plausible is by no means to be underestimated. Quite the contrary, it is to be welcomed that these physicists undertook the arduous task to adapt also more general philosophical notions to two physical ideas which have some very radical implications. However, the philosophical backing of physical ideas that emerged from these more general investigations has led to a situation that is by no means desirable. It has led to the belief in the uniqueness and the absolute validity of both of Bohr’s assumptions. It is, of course, admitted that the quantum theory may have to undergo some very decisive changes in order to cope with new phenomena […] it is also pointed out that, however large these changes may be, they will always leave untouched the two elements mentioned, viz. the indefiniteness of state descriptions and the relational character of the quantum-mechanical states, which, so it is added, cannot be replaced by different ideas without creating formal inconsistencies, or inconsistencies with experiment. - [1], page 313
We will discuss further some of these statements in relation to specific insights of PKF regarding Bohr’s views. Here we conclude this general discussion about the relation between physics and philosophy in the context of the foundations of QM with three remarks. First, in order to appreciate better PKF’s position above, it is useful to recall that PKF is writing in a historical period in which the debate about quantum foundations and interpretations had almost vanished, or at least it was not prominent among scientists anymore; a time in which the orthodox ”Copenhagen” interpretation, an instrumentalist or maybe simply a dismissive attitude toward any philosophical debate about physical theories, was entirely dominant in the physics community. Second, the problematic link between the risk of epistemological and scientific dogmatism and the use of ”no alternative arguments” in theory assessment and, more generally, the delicate interplay between scientific development and philosophy of physics, have recently been discussed in the context of the debate on non-empirical theory assessment in fundamental physics [9, 10, 11]. Finally, while PKF criticizes also Bohr as partially responsible for the dogmatic turn of the Copenhagen position, he was more generally full of praise, especially in later writings, for Bohr’s methodological and epistemological attitude, to be considered as providing a lesson about the proper interplay between physics and philosophy and for the need to avoid dogmatism.
@@
The idea of complementarity is therefore not just the result of having pursued a mistaken programme to the bitter end as Popper would want us to believe. Bohr, after all, did consider a purely statistical theory. He did consider such a theory despite the fact that it was not in line with his own point of view […] Bohr was originally interested in a generalization, not of classical statistical mechanics, but of classical particle mechanics. This part of Popper’s account is unobjectionable. However, Bohr was also open minded enough to generalize Einstein’s considerations and to develop and publish a purely statistical particle theory with waves acting as probability fields only. It was only after the refutation of this theory that he returned to his earlier philosophy - and this time with very good reasons. - [1], page 268
@@
(quoting Heisenberg) ’Bohr was primarily a philosopher, not a physicist, but he understood that natural philosophy . . . carries weight only if its every detail can be subjected to the . . . test of experiment’. His approach differed from that of the school-philosophers whom he regarded with a somewhat ’sceptical attitude, to say the least’ and whose lack of interest in ’the important viewpoint which had emerged during the development of atomic physics’ he noticed with regret. But it also differed, and to a considerable degree, from the spirit of what T. S. Kuhn has called a ’normal science’. Looking at Bohr’s method of research we see that technical problems, however remote, are always related to a philosophical point of view; they are never treated as ’tiny puzzles’ whose solution is valuable in itself, even if one has not the faintest idea what it means, and where it leads. […] Emphasis is put on matters of principle and minor discrepancies, or ’puzzles’ in the sense of Kuhn, instead of being deemphasized, and assimilated to the- older paradigm, are turned into fundamental difficulties by looking at them from a new direction, and by testing their background ’in its furthest consequences by exaggeration’. A noteworthy example of Bohr’s ’non-normal’ and rather metaphysical approach to physics is his scepticism in the face of the success of his own atomic model. […] Bohr not only upheld the approximate character of all microphysics, stationary states included, but he also stressed the ’formal’ character of the new laws, that is the fact that they aided in the prediction of phenomena, but did not yet provide any understanding (see below). His intention in those early days seems to have been the elaboration of a predictive scheme that could aid in the discovery of a physical theory in the full sense of the word, including an account even of the mechanism of the transitions. […] ’I noticed’, writes Heisenberg, ’that mathematical clarity had in itself no virtue for Bohr. He feared that the formal mathematical structure would obscure the physical core of the problem, and in any case he was convinced that a complete physical explanation should absolutely precede the mathematical formulation - [1], page 269-274

### v3

@@ 1 Introduction
In the area of foundations of physics, the analysis of technical aspects of physical theories and experimental results has the purpose of elucidating the conceptual pillars of the theories, which we use, in turn, to interpret what we see in experiments. In some cases, work on these foundations impact only our understanding of theories. Thus, it is valuable but not ground-breaking; it is interesting for scholars, but without much broader implications. In other cases, it shakes the whole edifice of theoretical physics, and has important implications for fundamental issues in philosophy of science and philosophy more generally. Thus, it is interesting for a broader set of philosophers, alongside physicists. This was certainly the case, and it still is [^1] It is also the case for the conceptual aspects of research on quantum gravity., for the foundations of quantum mechanics, whose puzzling aspects have provided exercises for prominent philosophers from its first developments.
@@
In the following, we focus on the contribution by P. K. Feyerabend (PKF) to the debate on the foundations and interpretation of quantum mechanics, on which he maintained a keen interest from early on but to which he contributed most actively in 1950s-1960s. The most relevant works are those collected in the first volume of PKF’s papers [1], to which we will refer throughout (rather than referring to the original publications).
@@
We will discuss several interesting points made by PFK. We group them in three categories. The first is about the general relation between physics and philosophy, which also serves as a framework for the philosophical analysis of quantum mechanics (and philosophy of physics at large). The second concerns the debate between realism and instrumentalism, applied to the interpretation of quantum mechanics, but also, again, as two broader perspectives on scientific theories. The third deals with specific insights on quantum mechanics itself, concerning the interpretation of quantum states, probabilities, Bohr’s view, and hidden variables.
@@
A result of our scan, beyond specific insights on quantum mechanics, is to illustrate how PFK constantly aimed to achieve a clearer understanding of the various issues he was facing, finding compelling rational arguments in favor of one particular viewpoint, while at the same time avoiding any form of dogmatism, even regarding the rational arguments he had just put forward, and regarding rationality itself (or the criteria adopted for evaluating his own arguments). This struggle characterizes his own version of rational analysis, dubbed ‘tightrope-walking rationality’ by Farrell [2], which we see in the context of the foundations of quantum mechanics and related philosophical issues. The dogmatic version of the two sides of any argument, e.g. realism vs instrumentalism, is what we see on the right and on the left of the tightrope we are forced to walk on, and we should avoid falling toward either side. The need to avoid dogmatism of any sort appears to be the strongest driver in PKF’s analysis, to the point where any perspective that is at risk of becoming a dogmatic position or facilitating a dogmatic attitude is immediately objected to exactly on that basis. Thus, avoiding dogmatism turns quickly from a normative methodological rule for philosophical inquiry into an evaluative criterion of philosophical positions in itself. We will emphasize various concrete instances of application of such tightrope-walking rationality in what follows.
@@
Before we start, it is worth emphasizing what this contribution is about and what is not aiming at. It is not intended as a scholarly contribution about PKF’s work and thought. We do not attempt a precise or detailed analysis of what PKF really meant or of how PKF’s thinking evolved over the years. We do not frame PKF’s thinking within the historical debate on quantum foundations, and we do not assess his actual impact on the same debate. For proper scholarly work on PKF, especially in relation to contemporary debates in philosophy of science (e.g. on realism, perspectivism, pluralism, scientific methodology, etc), see for example [3, 4, 5] or the cited [2], and, in broader terms [6, 7]. For further analysis on PKF’s philosophy of physics, see instead [8, 9, 10, 11, 12, 13]. What we aim to do is much more limited. Being interested ourselves in the many puzzles of quantum foundations, we scan PKF’s writings for useful insights that could help us improving our own understanding of them, and could suggest ways to make technical or conceptual progress. In particular, we ask PKF for advice about how one should seek some form of scientific realism behind quantum mechanics, about the complex relation between a scientific theory and reality, and about how to be critical toward our own attitudes when presented with a new set of phenomena and a new candidate description of them. Also in this respect, our contribution should be seen as a modest addition to the scholarly work on this topic, examples being [8, 14] It will be, then, an exercise in picking up useful inputs for current research, and suggestive stimuli to be put to fruition at a second stage. We will try, of course, to be careful in our reading and understanding PKF’s thoughts, and to avoid misinterpretations or cherry-picking or other forms of unfaithfulness to PKF’s intentions and ideas. In case we fail despite our best efforts, and some cherry-picking turns out to be happening, even if unintended, we hope that at least the cherries are good.
@@ 2 Relation between physics and philosophy, in the foundations of QM
An immediate point is PKF’s preference for philosophical analyses that, when not directly done by physicists, it is proceeding in parallel with scientific developments and to some extent guided by them; this point has been emphasized already in [9, 15]. On the one hand, an important goal of the philosophy of physics is, in PKF’s view, to contribute to the scientific progress; on the other hand, a physical argument or result that supports or suggests a certain philosophical perspective becomes the best argument in its favor and has more weight than abstract philosophical reasoning going in the same or opposite direction. This applies to metaphysical issues, as well as epistemological or methodological ones. In this sense, one can understand the many occasions in which PKF expresses discomfort or outright annoyance with philosophical positions that appear oblivious to the hard-won results by scientists or that are based on a superficial picture of scientific theories, and of how they have been obtained and tested, or with methodological rules that have no relation with the practice of science.
We would suggest that a way to understand the spirit of PKF’s attitude in this regard is to consider his tightrope-walking rationality not so much as an abstract theory of rationality to be applied (or imposed) onto science and scientists, but, quite the contrary, as a model of rationality deduced or abstracted from the way actual scientists work, from the necessarily fallibilist and wavering proceeding of real-world science.
In a general survey on scientific realism in physics, for example, PKF praises the ideas of Maxwell and Mach because “They are also more subtle. They were developed in close connection with research […]. But one feels a sense of relief when transferred from the fruitless technicalities and ontological primitivisms of modern ’philosophers’ to the brief, simple, but profound remarks of these scientists”. [[1], page 11].
@@
Lakatos says I use ‘some inconsistencies and waverings in Bohr’s position for a crude apologetic falsification of Bohr’s philosophy’. This is the language of the rationalist who looks at research from the safe distance of his office and wants to run science according to one method and base it on one philosophy and who fails to understand that researchers adapt their ideas to the always changing research situation. Bohr does not ‘waver’because he does not know what to do, but because new discoveries are constantly being made and because he takes them into account. - [1], page 294, footnote 100
@@
Ultimately, physical arguments (and results) trump philosophical prejudices, be they methodological or ontological. This is as it should be, to make science progress. On the methodological side, this attitude is also at the basis of PKF’s methodological anarchism [16], since it allows for different methodological prescriptions to be employed in different circumstances, and any given methodological rule to be violated, if deemed useful to achieve scientific progress (a similar understanding of PKF’s methodological anarchism has been advocated in [17]).
@@
Does this mean that philosophy plays no role in scientific progress? Obviously not. Philosophical arguments can guide or support physics; a certain philosophical perspective can prove fertile in driving scientific progress or solving conceptual puzzles encountered by scientists. Ontological preferences are means to physical understanding; they are either starting points for new scientific constructions or results of (analysis of) existing ones. Philosophical arguments can also be a good motivation to oppose a scientific theory or viewpoint, but only as starting point to develop alternative scientific theories. It is at the level of physics that the debate will have to be settled, with the physical reasons being given appropriate (and greater) weight. Philosophical arguments can win a philosophical battle, never one about physics. This implies, in particular, that philosophical as well as physical arguments should be considered in relation to specific theories, those that we are scrutinizing for viability, and not in general, abstract terms.
@@
PKF is very clear on this point, when discussing the debate around the heliocentric view and then using it to elucidate the parallel debate on the realist vs instrumentalist interpretation of quantum theory (see for example [1], page 185-186). In [18], PKF’s position regarding the ‘realism vs instrumentalism’debate in discussed from a similar understanding of his attitude.
@@
This is a warning that should be raised also in front of physicists trying to wrap their theories in fancy philosophical clothes, trying to make their insights look deeper, or trying to call philosophical frameworks in further support of their physical theories. This move can have value but one has to be very careful. Bad philosophical arguments can end up obscuring good physical motivations and bring unnecessary trouble to a physical theory or viewpoint that could have stood on more solid grounds without. This is what has happened, according to PKF, to the ”instrumentalist”  interpretation of quantum mechanics. In the attempt to further justify such position beyond the physical arguments by linking it with more general positivism, quantum physicists ended up offering easy points of attack to their opponents. Attacking their positivist position and refuting it with good philosophical arguments made their view of quantum theory appear weaker, despite their being in possession of strong physical results supporting it. So strong, in fact, that they did not need to be further justified in terms of any positivist philosophy. In the same spirit, PKF warns about expressing or interpreting physical theories in a language adapted to or coming from a specific philosophical perspective. In the specific case of quantum transitions, for example, PKF insists that the absence of intermediate states is something to be argued for because of the empirical evidence and of the content of quantum theory that successfully accounts for it, while arguing for this conclusion or justifying the theory on the basis of a philosophical doctrine (here, positivism) does not help:
@@
[…] the quantum postulate does not merely exclude the knowledge of, or the observability of, the intermediate states; it excludes these intermediate states themselves. Nor must the argument be read as asserting, as is implied in many presentations by positivist-minded physicists, that the intermediate states do not exist because they cannot be observed. For it refers to a postulate (the quantum postulate) which deals with existence, and not with observability. […] Physicists who have adopted the positivistic principle that things which cannot be observed do not exist try to justify the indefiniteness of state descriptions by a combined reference to the fact that they cannot be observed and to this principle. Philosophers immediately expose the fallacy of the argument (if they are anti-positivists, that is) and think that they have thereby shown the existence, or at least the physical possibility, of sharp states. This is, of course, not correct […][1], page 187-190
@@
[…] modern physics has found new physical reasons why its own most important theory, the quantum theory, cannot be anything but an instrument of prediction. These reasons are of precisely the same character as were Ptolemy’s: a realistic interpretation of the quantum theory is bound to lead to incorrect predictions. Admittedly, in the arguments usually presented these physical difficulties are almost buried beneath an unacceptable positivism. However, this does not mean that they do not exist […] - [1], page 187-190
@@
the instrumentalism in the quantum theory is not a purely philosophical affair that can be disputed away by general arguments in favour of realism. A realistic alternative to the idea of complementarity is likely to be successful only if it implies that certain experimental results are not strictly valid. It therefore demands the construction of a new theory, as well as demonstration that this new theory is experimentally at least as valuable as the theory that is being used at present. This is a formidable task indeed, and a task that is not even recognized by the purely philosophical champions of realism in microphysics. - [1], page 193-194
@@
Bohm’s physical ideas are original, refreshing, and sorely needed in a time of complacency with respect to fundamentals. But the philosophical standpoint taken up with respect to both physics and cosmology is traditional, and perhaps even reactionary - [1], page 220
In the same discussion, PKF introduces another important element: a sound position can quickly become counterproductive and represent an obstacle to further progress the very moment it becomes dogmatic, no matter how many solid arguments support it. A position that is well supported by physical arguments and observations, that is embodied in the best physical framework we were able to find, and that it is proving fertile for further scientific progress becomes dogmatic when it is declared to be necessarily part of any future theory, to stand without possible alternatives, rather than merely current ones, as something having absolute validity. Also here one can see the ideal tightrope-walking philosopher or scientist moving the balance quickly away from the full (and possibly uncritical) embracing of a theory and towards a more skeptical or simply actively critical attitude, to avoid falling into the pit of dogmatism. This dogmatic turn is what PKF attributes to the ”orthodox” Copenhagen interpretation and to the physicists holding it. In his words:
The idea of complementarity can be interpreted […]as an attempt to provide an intuitive picture for an existing theory, viz. wave mechanics, and as a heuristic principle guiding future research. This interpretation is undogmatic as it admits the possibility of alternatives, and even of preferable alternatives. A physicist who looks at complementarity in this way will regard it as an interesting fact about quantum theory that it is compatible with a relational point of view where interaction is a necessary condition of the meaningful applicability of terms which within classical physics (relativity included) are definable without such reference. He will also point out that there exist no satisfactory alternatives. But he will never go as far as to assert that such alternatives will never be found, or that they would be logically inconsistent, or that they would contradict the facts. But Bohr’s idea of complementarity can also be interpreted […] as a basic philosophical principle which is incapable of refutation and to which any future theory must conform. Bohr himself most certainly took this stronger point of view. - [1], page 221-222
@@
it has been attempted, both by Bohr, and by some other members of the Copenhagen circle, to give greater credibility to these ideas by incorporating them into a whole philosophical (ontological) system that comprises physics, biology, psychology, sociology and perhaps even ethics. Now the attempt to relate physical ideas to a more general background and the correlated attempt to make them intuitively plausible is by no means to be underestimated. Quite the contrary, it is to be welcomed […] However, the philosophical backing of physical ideas that emerged from these more general investigations has led to a situation that is by no means desirable. It has led to the belief in the uniqueness and the absolute validity of both of Bohr’s assumptions […] viz. the indefiniteness of state descriptions and the relational character of the quantum-mechanical states, which, so it is added, cannot be replaced by different ideas without creating formal inconsistencies, or inconsistencies with experiment. - [1], page 313
We will further discuss some of these statements in relation to specific insights of PKF regarding Bohr’s views. Here, we conclude this general discussion with three remarks. First, to better appreciate PKF’s position above, it is useful to recall that PKF is writing in a historical period in which the debate about quantum foundations and interpretations had almost vanished, or at least it was not prominent among scientists anymore; a time in which the orthodox “Copenhagen”interpretation, an instrumentalist or maybe simply a dismissive attitude toward any philosophical debate about physical theories, was entirely dominant in the physics community. Second, the problematic link between the risk of epistemological and scientific dogmatism and the use of “no alternative arguments”in theory assessment and, more generally, the delicate interplay between scientific development and philosophy of physics, have recently been discussed in the context of the debate on non-empirical theory assessment in fundamental physics [19, 20, 21]. Finally, while PKF criticizes Bohr as partially responsible for the dogmatic turn of the Copenhagen position, he was more generally full of praise, especially in later writings, for Bohr’s methodological and epistemological attitude, as providing a lesson about the proper interplay between physics and philosophy and for the need to avoid dogmatism.
@@
The idea of complementarity is therefore not just the result of having pursued a mistaken programme to the bitter end as Popper would want us to believe. Bohr, after all, did consider a purely statistical theory. He did consider such a theory despite the fact that it was not in line with his own point of view […] It was only after the refutation of this theory that he returned to his earlier philosophy - and this time with very good reasons. - [1], page 268
@@
(quoting Heisenberg) ’Bohr was primarily a philosopher, not a physicist, but he understood that natural philosophy . . . carries weight only if its every detail can be subjected to the . . . test of experiment’. His approach differed from that of the school-philosophers whom he regarded with a somewhat ’sceptical attitude, to say the least’ and whose lack of interest in ’the important viewpoint which had emerged during the development of atomic physics’ he noticed with regret. But it also differed, and to a considerable degree, from the spirit of what T. S. Kuhn has called a ’normal science’. Looking at Bohr’s method of research we see that technical problems, however remote, are always related to a philosophical point of view; they are never treated as ’tiny puzzles’ whose solution is valuable in itself, even if one has not the faintest idea what it means, and where it leads. […] Emphasis is put on matters of principle and minor discrepancies, or ’puzzles’ in the sense of Kuhn, instead of being deemphasized, and assimilated to the- older paradigm, are turned into fundamental difficulties by looking at them from a new direction, and by testing their background ’in its furthest consequences by exaggeration’. - [1], page 269-274